I'm really stuck on the below question and I've been coming back to it with different ways to complete it but just can't wrap my head around it. Many thanks.
John is one year older than Pat. Mary is twice as old as John. The sum of their age is 63. How old is John.

Found 2 solutions by Theo, MathTherapy:
Answer by Theo(13342)

let j = the age of john
let p = the age of pat
let m = the age of mary
let s = the sum of their ages
then s = j + p + m

john is one year older than pat.
this get you j = p + 1
solve for p to get p = j - 1

mary is twice the age of john.
this gets you m = 2 * j

the sum of their ages is 63.
this gets you s = 63.
since s = j + p + m, then j + p + m = 63.
since p = j - 1 and m = 2 * j, then j + p + m = 63 becomes:
j + j - 1 + 2 * j = 63
add 1 to both sides of this equation to get:
j + j + 2 * j = 64
combine like terms to get:
4 * j = 64
solve for j to get:
j = 16

since p = j - 1, then p = 15
since m = 2 * j, then m = 32

you get:
j = 16
p = 15
m = 32

s = j + p + m becomes sum s = 16 + 15 + 32 which becomes s = 63
j = p + 1 becomes j = 15 + 1 which becomes j = 16
m = 2 * j becomes m = 2 * 16 which becomes m = 32

john is 16
pat is 15
mary is 32

Answer by MathTherapy(10552)


I'm really stuck on the below question and I've been coming back to it with different ways to complete it but just can't wrap my head around it. Many thanks.
John is one year older than Pat. Mary is twice as old as John. The sum of their age is 63. How old is John.
Let John's age be J
Then Pat's is: J - 1
Also, Mary's is: 2J
With their ages summing to 63, we get: J + J - 1 + 2J = 63
4J = 64
J, or